OpenStudy (anonymous):

The equation of line CD is y = −2x − 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5).

OpenStudy (anonymous):

y=-2x has to start like this to be parallel (same slope) Can you figure out the rest of the equation?

OpenStudy (anonymous):

wats sup

OpenStudy (anonymous):

Not much. At work... boring rainy day

OpenStudy (anonymous):

same here... ummm i believe you have to replace the y with 4, so 4=2x

OpenStudy (anonymous):

so im thinking divide 4/-2, which equals -2

OpenStudy (anonymous):

(4, 5). this point has to satisfy the equation x is first, y second

OpenStudy (anonymous):

ohhh yea, so do you divide the 5/-2?

OpenStudy (anonymous):

y = −2x + or - something will be the solution at (4, 5) 5=-2*4 5=-8 so the - or + something is??

OpenStudy (anonymous):

-3 must equal b then

OpenStudy (anonymous):

so the final answer could be y=-2x-3

OpenStudy (anonymous):

5=-8 -3 =-11 Still nope 5=-8 +-?

OpenStudy (anonymous):

5=-8 +-? Instead of +-? you can write lets say z 5=-8 +z z=?

OpenStudy (anonymous):

where did the -11 from

OpenStudy (anonymous):

im sorry come from

OpenStudy (anonymous):

You wrote -3 -8-3=-11 that is not 5

OpenStudy (anonymous):

y=-2x-3 this line does not contain (4,5) as 5=-8-3=-11 not cool

OpenStudy (anonymous):

ok, so what i did was take the 5 and subtract 8 from it, which is -3=b since you solve for b and i was just thinking that you could not change the slope

OpenStudy (anonymous):

i honestly dont know where to go with this

OpenStudy (anonymous):

5=-8 + b Lets use b than :) Solve the above!

OpenStudy (anonymous):

i got -27=b

OpenStudy (anonymous):

because you plug in the 4 right?

OpenStudy (anonymous):

No :( Sorry

OpenStudy (anonymous):

5=-8 + b Add 8 to both sides 5+8=-8+8+b 13=b

OpenStudy (anonymous):

wow, i found my mistake now

OpenStudy (anonymous):

i had to add 8 to both sides. Thanks man, sorry for my incompetence

OpenStudy (anonymous):

Oh that is nice :) It is always the best way to learn something

OpenStudy (anonymous):

No probs :) In a few years you will get there